### v2

# FE-LWS: Refined Image-Text Representations via Decoder Stacking and Fused Encodings for Remote Sensing Image Captioning

Abstract
Remote sensing image captioning (RSIC) aims to generate meaningful textual descriptions from remote sensing images using encoder-decoder frameworks. Although previous efforts have focused primarily on improving the design of decoders, significant challenges persist in capturing the complex spatial and semantic characteristics of remote sensing data. In addition, reliance on a single encoder or decoder often limits the representational capacity of the model. To address these limitations, we propose a novel framework that fuses features from two complementary CNN encoders and integrates a stacked GRU decoder with weighted averaging to better model multilevel semantics. Experimental evaluations demonstrate that the proposed approach outperforms existing baselines, delivering more accurate and contextually rich captions for remote sensing images.
Index Terms:
Remote Sensing Image Captioning (RSIC), Encoder-Decoder model, Convolutional Neural Network (CNN), Gated Recurrent Unit (GRU), Fusion-based Encoders (FE), Weighted-based Stacking (WS).
I Introduction
Remote Sensing Image Captioning (RSIC) is a task in the machine learning domain that focuses on extracting meaningful information from a remote sensing images. The encoder-decoder framework is the dominant approach in RSIC, where a CNN-based encoder (such as ResNet [1], ConvNext [2]) extracts image features and a decoder (e.g., LSTM [3, 4], GRU [5], or SVM [6]) generates captions in sequence-to-sequence manner.
Although decoder optimization has received considerable attention [3, 7], the encoder’s role in capturing spatial and semantic diversity remains underexplored. Most of the methods are based on a single CNN, which limits the representational richness. To address this, we propose a fusion-based encoder that integrates features from two complementary CNNs, enabling a more detailed and robust image understanding.
Similarly, most existing models employ a single decoder, which restricts the modeling of hierarchical semantics and long-range dependencies [4, 8, 9]. We adopt a stacked GRU-based decoder [10, 11] to sequentially refine features across layers, improve contextual representation, and mitigate vanishing gradients. To avoid losing intermediate information, we aggregated the outputs of all decoder layers using a weighted averaging strategy, allowing the model to capture both low and high-level semantics more effectively.
The major contributions of this letter are as follows.
  • A fusion-based encoding strategy is proposed to enhance image representation and understanding.
  • A novel multi-layer stacked GRU decoder with weighted averaging is proposed to improve sequence modeling.
  • The proposed model is thoroughly evaluated against state-of-the-art methods, demonstrating its notable performance improvement.
The structure of the letter is as follows: Section II introduces our proposed approach, focusing on fusion-based encoding and stacking-based decoding techniques. Section III describes the experimental setup, covering datasets, model configurations, and evaluation metrics. Table IX presents the results of various aspects, including ablation studies of our model, and provides a comparative analysis with baseline models through both quantitative and qualitative evaluations, supported by visual examples. Finally, Section V summarizes the key findings of the letter and highlights potential avenues for future research.
II Proposed Method
Figure 1: Architecture of the Proposed Model
The overall architecture is illustrated in Figure 1. We adopt an encoder-decoder framework, where the encoder ($F$) consists of two distinct CNNs whose outputs are concatenated and processed through a linear transformation layer ($L_{1}$). For decoding, a tokenized input sequence passes through an embedding layer ($E$) and then into a stack of Gated Recurrent Units (GRU) ($G$) to model temporal dependencies. The outputs from $L_{1}$ and $G$ are concatenated and passed through two additional linear layers ($L_{2}$ and $L_{3}$), with $L_{3}$ serving as a softmax layer to generate word probabilities.
II-A Fusion of Different Encoders
Figure 2: Architecture of the Fusion Technique Used in Our Model
The proposed method integrates features extracted from two widely used CNN architectures. The fusion strategy is illustrated in Figure 2. Given the vast number of CNN models available, an exhaustive comparison of all individual models and their pairwise combinations is impractical. Therefore, this study focuses on four prominent architectures: ConvNext [2], ResNet [1], ResNext [12], and DenseNet [13]. These models have demonstrated strong feature extraction capabilities in RSIC [4, 14].
TABLE I: Performance Comparison of Individual and Fusion-Based CNN Encoders on the SYDNEY Dataset
CNN BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
ConvNext 0.7531 0.6645 0.5982 0.5376 0.3750 0.6951 2.3954
ResNet 0.7363 0.6502 0.5785 0.5138 0.3621 0.6809 2.2526
ResNext 0.7415 0.6556 0.5845 0.5209 0.3625 0.6706 2.2912
DenseNet 0.7214 0.6331 0.5621 0.4935 0.3419 0.6314 2.1444
[…]
ConvNext-ResNext 0.7582 0.6727 0.6039 0.5438 0.3868 0.7164 2.3882
ConvNext-DenseNet 0.7419 0.6571 0.5823 0.5264 0.3822 0.6911 2.3755
ResNet-ResNext 0.7517 0.6664 0.5924 0.5243 0.3789 0.6963 2.3515
ResNet-DenseNet 0.7246 0.6413 0.5722 0.5101 0.3644 0.6798 2.2488
ResNext-DenseNet 0.7322 0.6514 0.5814 0.5299 0.3714 0.6833 2.2655
TABLE II: Performance Comparison of Individual and Fusion-Based CNN Encoders on the UCM Dataset
CNN BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
ConvNext 0.8369 0.7804 0.7312 0.6853 0.4689 0.8037 3.3990
ResNet 0.8297 0.7705 0.7178 0.6683 0.4506 0.7903 3.3329
ResNext 0.8170 0.7607 0.7118 0.6658 0.4583 0.7736 3.3165
DenseNet 0.8222 0.7648 0.7211 0.6714 0.4674 0.7833 3.3419
[…]
ConvNext-ResNext 0.8406 0.7796 0.7289 0.6813 0.4656 0.7996 3.4132
ConvNext-DenseNet 0.8416 0.7834 0.7319 0.6877 0.4798 0.8101 3.4522
ResNet-ResNext 0.8334 0.7739 0.7236 0.6723 0.4580 0.7944 3.4108
ResNet-DenseNet 0.8311 0.7825 0.7230 0.6711 0.4596 0.7998 3.3712
ResNext-DenseNet 0.8367 0.7898 0.7292 0.6736 0.4674 0.8112 3.3919
TABLE III: Performance Comparison of Individual and Fusion-Based CNN Encoders on the RSICD Dataset
CNN BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
ConvNext 0.6327 0.4636 0.3602 0.2903 0.2487 0.4648 0.8272
ResNet 0.6322 0.4621 0.3557 0.2835 0.2489 0.4649 0.8206
ResNext 0.6246 0.4525 0.3463 0.2639 0.2294 0.4448 0.8117
DenseNet 0.6298 0.4566 0.3504 0.2668 0.2207 0.4519 0.8167
[…]
ConvNext-ResNext 0.6414 0.4717 0.3676 0.2970 0.2551 0.4741 0.8429
ConvNext-DenseNet 0.6392 0.4634 0.3699 0.3001 0.2588 0.4744 0.8392
ResNet-ResNext 0.6427 0.4734 0.3684 0.2974 0.2558 0.4743 0.8431
ResNet-DenseNet 0.6345 0.4578 0.3542 0.2741 0.2443 0.4694 0.8346
ResNext-DenseNet 0.6282 0.4552 0.3519 0.2712 0.2311 0.4560 0.8211
As shown in Tables III, III and III, the combination of features from two different CNNs improves performance compared to the use of a single CNN. This fusion does not add significant computational cost, as CNNs are used solely for feature extraction and remain frozen during training. Since the encoder parameters are not updated, the training remains efficient while benefiting from the diverse spatial and semantic information captured by multiple CNNs. Among the combinations, ConvNext and ResNet yield the best results due to their complementary properties. ConvNext effectively captures global spatial patterns, while ResNet retains fine-grained details through residual connections. Previous studies [4, 14] also highlight their superiority in RSIC tasks. Their fusion results in richer and more robust feature representations, making ConvNext and ResNet the chosen pair for all subsequent experiments in this study.
II-B Stacking of Multiple Decoders
Our model uses the Gated Recurrent Unit (GRU) [15] as the decoder. The GRU effectively manages the flow of information through its gating mechanisms, allowing it to maintain crucial temporal dependencies while minimizing computational complexity. Conventional encoder-decoder architectures typically rely on a single decoder to generate captions in a sequence-to-sequence manner [16, 7, 6, 4]. However, recent work has shown that stacking multiple decoders enhances contextual modeling and hierarchical learning by refining intermediate representations and improving long-term dependency capture [17, 18, 19]. In our approach, we employ a stacked GRU structure in the decoder to enhance the model’s ability to process sequential data and to improve its robustness in RSIC. However, simply stacking GRUs sequentially and relying solely on the final layer’s output does not always yield the optimal result because deeper networks can suffer from vanishing-gradient problems [20, 21], and important intermediate features might be lost if only the last GRU’s output is considered. Furthermore, the information flow between stacking layers may not be efficiently balanced, leading to suboptimal learning. To address these challenges, we explore four different stacking strategies.
TABLE IV: Stacked Decoder Results on SYDNEY Dataset
DC Stack BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
1 NS 0.7679 0.6814 0.6084 0.5458 0.3801 0.7263 2.4515
[…]
CS 0.7784 0.7012 0.6400 0.5853 0.4287 0.7319 2.5614
GWS 0.7856 0.7106 0.6482 0.5923 0.4195 0.7286 2.5691
LWS 0.8096 0.7397 0.6580 0.6114 0.4369 0.7461 2.6127
[…]
CS 0.7680 0.6901 0.6247 0.5654 0.4088 0.7139 2.4645
GWS 0.8049 0.7319 0.6696 0.6198 0.4348 0.7479 2.6658
LWS 0.8349 0.7589 0.6952 0.6369 0.4554 0.7767 2.9230
[…]
CS 0.7576 0.6758 0.6099 0.5520 0.3993 0.7017 2.3367
GWS 0.7953 0.7212 0.6640 0.6134 0.4266 0.7403 2.6238
LWS 0.8144 0.7423 0.6849 0.6338 0.4466 0.7587 2.7044
[…]
CS 0.7646 0.6769 0.6136 0.5577 0.4039 0.7136 2.4011
GWS 0.7998 0.7282 0.6668 0.6111 0.4288 0.7429 2.5669
LWS 0.8276 0.7537 0.6904 0.6306 0.4509 0.7706 2.7253
TABLE V: Stacked Decoder Results on UCM Dataset
DC Stack BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
1 NS 0.8457 0.7919 0.7533 0.7166 0.4833 0.8163 3.4255
[…]
CS 0.8416 0.7860 0.7379 0.6910 0.4642 0.8038 3.5311
GWS 0.8572 0.8012 0.7510 0.7028 0.4745 0.8089 3.4716
LWS 0.8663 0.8157 0.7711 0.7276 0.4948 0.8307 3.5485
[…]
CS 0.8398 0.7848 0.7356 0.6881 0.4622 0.8019 3.5216
GWS 0.8678 0.8158 0.7706 0.7266 0.4889 0.8265 3.6196
LWS 0.8805 0.8305 0.7862 0.7511 0.5014 0.8497 3.6806
[…]
CS 0.8397 0.7844 0.7374 0.6897 0.4685 0.8072 3.4939
GWS 0.8612 0.8058 0.7581 0.7116 0.4769 0.8173 3.5104
LWS 0.8661 0.8090 0.7614 0.7170 0.4849 0.8215 3.5838
[…]
CS 0.8361 0.7818 0.7376 0.6915 0.4633 0.7958 3.4776
GWS 0.8468 0.7904 0.7416 0.6952 0.4806 0.8143 3.4959
LWS 0.8613 0.8054 0.7593 0.7157 0.4896 0.8147 3.6018
TABLE VI: Stacked Decoder Results on RSICD Dataset
DC Stack BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
1 NS 0.6452 0.4767 0.3732 0.3087 0.2609 0.4827 0.8594
[…]
CS 0.6536 0.4871 0.3846 0.3146 0.2698 0.4888 0.8641
GWS 0.6586 0.4921 0.3881 0.3166 0.2712 0.4921 0.8822
LWS 0.6683 0.5007 0.3969 0.3265 0.2836 0.5027 0.8818
[…]
CS 0.6531 0.4869 0.3818 0.3116 0.2671 0.4894 0.8697
GWS 0.6614 0.5014 0.3984 0.3314 0.2911 0.5134 0.9196
LWS 0.6796 0.5124 0.4073 0.3365 0.2944 0.5248 0.9423
[…]
CS 0.6463 0.4797 0.3757 0.3043 0.2530 0.4698 0.8441
GWS 0.6511 0.4846 0.3806 0.3091 0.2637 0.4846 0.8747
LWS 0.6596 0.4924 0.3873 0.3165 0.2744 0.4938 0.8823
[…]
CS 0.6425 0.4734 0.3726 0.3014 0.2490 0.4636 0.8361
GWS 0.6502 0.4812 0.3745 0.2894 0.2511 0.4511 0.8426
LWS 0.6519 0.4843 0.3800 0.3089 0.2692 0.4886 0.8653
II-B1 Simple Stacking
The traditional stacking approach, called Simple Stacking (SS), directly utilizes the output of the final decoder layer as the final output of the system. The overall output of the SS system is given below.
$O=D_{n}$ (1)
II-B2 Concatenation-based Stacking
In Concatenation-based Stacking (CS), the outputs of all decoder layers are concatenated to form the final output, thus preserving a richer set of feature representations. Then, the overall output $O$ of the CS system is given below.
$O=D_{1}|D_{2}|...|D_{n}$ (2)
II-B3 Global Weighted-based Stacking
In Global Weighted-based Stacking (GWS), a weighted average of the output of all decoder layers is calculated using a single global weight, ensuring a balanced contribution from each layer. The output of the GWS system is given below.
$O=\sum_{i=1}^{n}\alpha_{i}{D_{i}}$ (3)
II-B4 Local Weighted-based Stacking
In Local Weighted-based Stacking (LWS), a weighted average of the decoder output is calculated using element-specific weights, allowing for finer control over the contribution of each feature to the final output. The output of the LWS system is given below.
$O=\sum_{i=1}^{n}W_{i}\odot D_{i}$ (4)
Here, $n$ denotes the number of decoders, and $D_{i}$ represents the output of the $i^{\text{th}}$ decoder, while $O$ is the final output of the stacked system. The symbol |||| indicates the concatenation operation. The parameter $\alpha_{i}$ represents a scalar weight, while $W_{i}$ is a vector weight of the same size as $D_{i}$ for each $D_{i}$. Furthermore, $0\leq\alpha_{i}\leq 1$ with constraint $\sum_{i=1}^{n}\alpha_{i}=1$, and $0\leq W_{i}^{j}\leq 1$ with $\sum_{i=1}^{n}W_{i}^{j}=1$, where $W_{i}^{j}$ denotes the value of the characteristic $j^{\text{th}}$ of $W_{i}$. Both $\alpha=\{\alpha_{i}\mid i=1,\dots,n\}$ and $W=\{W_{i}\mid i=1,\dots,n\}$ are trainable parameters.
The impact of the depth of the decoder is evaluated in Tables VI, VI and VI, where the number of stacked GRUs (DC) is treated as a hyperparameter ranging from two to five. Configurations beyond five are excluded due to increased computational cost and diminishing returns.
In the SYDNEY and UCM datasets( Tables VI and VI), deeper stacks under SS and CS show inconsistent trends, often degrading beyond two layers. SS, in particular, underperforms the single-decoder baseline (NS), probably due to vanishing gradients and limited training data [22]. RSICD maintains stable performance up to three layers, with declines beyond that point, indicating sensitivity to the depth of the decoder.
From the results, we observe that among the stacking strategies, SS and CS are less effective. SS eliminates intermediate outputs, while CS introduces redundancy by treating all outputs equally. Weighted strategies offer a better balance: GWS uses global weights but lacks temporal flexibility, whereas LWS applies adaptive element-wise weights, enabling finer control over temporal dependencies. LWS consistently achieves the best results, with three stacked decoders yielding optimal performance across all datasets.
III Experimental Setup
We use a conventional encoder-decoder architecture. Each GRU on the decoder stack ($G$) has an output size of $256$. Linear layers $L1$, $L2$, and $L3$ have output sizes of $256$, $512$, and one more than the vocabulary size, respectively. GELU is used as the activation function for $L1$ and $L2$, while $L3$ uses SoftMax. A dropout rate of $0.5$ is applied before $L1$, after the embedding layer ($E$), before the sequence loss layer ($SL$), and between consecutive GRUs to prevent overfitting. The embedding size is set to $256$.
Training is carried out with a batch size of $64$ using the Adam optimizer (learning rate $10^{-4}$) and categorical cross entropy as a loss function. An early stopping criterion with a patience of eight epochs is used, monitoring ROUGE-L on the validation set. For caption generation, we apply the comparison-based beam search [3, 7] with a beam width of five, using four similar images. The comparison score is calculated as the arithmetic mean of BLEU-2, METEOR, and ROUGE-L.
The experiments were carried out using TensorFlow within a Docker environment with GPU support, on a host machine equipped with a NVIDIA RTX A6000 GPU with $50GB$ memory and $132GB$ RAM. During manuscript preparation. ChatGPT and Grammarly were used solely for rephrasing and language refinement, without generating any original technical content.
III-A Datasets Used
The performance of the model was evaluated using three widely recognized RSIC datasets: SYDNEY [16], UCM [16], and RSICD [22]. The details of these datasets are provided in the supplementary material. In this study, we used the corrected versions of these datasets [3], which address common issues such as spelling mistakes, grammatical errors, and inconsistent dialect variations of English. We used the same train-validation-test split as defined in these datasets.
III-B Performance metrics
The evaluation metrics used in our model include BLEU [23], METEOR [24], ROUGE [25], and CIDEr [26], which are widely adopted in RSIC tasks [16, 22, 3, 4]. More details are provided in the supplementary material.
IV Results
TABLE VII: Performance Comparison of Model Variants Through Ablation Studies on SYDNEY Dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
C-NS 0.7531 0.6645 0.5982 0.5376 0.3750 0.6951 2.3954
R-NS 0.7363 0.6502 0.5785 0.5138 0.3621 0.6809 2.2526
FE-NS 0.7679 0.6814 0.6084 0.5458 0.3801 0.7263 2.4515
C-LWS 0.8141 0.7321 0.6610 0.6054 0.4369 0.7584 2.7472
R-LWS 0.8056 0.7211 0.6527 0.5924 0.4216 0.7364 2.5743
FE-LWS 0.8349 0.7589 0.6952 0.6369 0.4554 0.7767 2.9230
TABLE VIII: Performance Comparison of Model Variants Through Ablation Studies on UCM dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
C-NS 0.8369 0.7804 0.7312 0.6853 0.4689 0.8037 3.3990
R-NS 0.8297 0.7705 0.7178 0.6683 0.4506 0.7903 3.3329
FE-NS 0.8457 0.7919 0.7533 0.7166 0.4833 0.8163 3.4255
C-LWS 0.8655 0.8121 0.7644 0.7192 0.5086 0.8367 3.4958
R-LWS 0.8614 0.8034 0.7526 0.7049 0.4910 0.8194 3.4621
FE-LWS 0.8805 0.8305 0.7862 0.7511 0.5014 0.8497 3.6806
TABLE IX: Performance Comparison of Model Variants Through Ablation Studies on RSICD dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
C-NS 0.6327 0.4636 0.3602 0.2903 0.2487 0.4648 0.8272
R-NS 0.6322 0.4621 0.3557 0.2835 0.2489 0.4649 0.8206
FE-NS 0.6452 0.4767 0.3732 0.3087 0.2609 0.4827 0.8594
C-LWS 0.6671 0.5008 0.3996 0.3302 0.2843 0.5066 0.8954
R-LWS 0.6645 0.4981 0.3963 0.3257 0.2849 0.5010 0.8684
FE-LWS 0.6796 0.5124 0.4073 0.3365 0.2944 0.5248 0.9423
This section presents a comprehensive evaluation of the proposed model through ablation studies, quantitative comparisons with baseline methods, and visual analysis. The goal is to assess the individual and combined contributions of encoder fusion, decoder stacking, and caption selection strategy.
IV-A Ablation Studies
To evaluate the contribution of individual components, we compare four encoder-decoder configurations: single encoder with single decoder, fusion-based encoder with single decoder, single encoder with LWS decoder and the proposed fusion-based encoder with LWS decoder.
The results, shown in Tables IX, IX and IX, follow the naming convention where C denotes ConvNext,R denotes ResNet, andFE indicates their fusion. For decoders,NS refers to a single decoder andLWS to Local Weighted-based Stacking with three GRUs (see Section II-B).
The results indicate that both encoder fusion and decoder stacking improve performance, with the FE-LWS configuration achieving the highest scores across all datasets. In particular, LWS offers greater performance gains than encoder fusion alone, highlighting its effectiveness even when used with a single CNN.
IV-B Numerical Evaluation
TABLE X: Experimental Results of Baseline Methods and Different Searching Techniques on SYDNEY Dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
R-BOW [22] 0.5310 0.4076 0.3319 0.2788 0.2490 0.4922 0.7019
L-FV [22] 0.6331 0.5332 0.4735 0.4303 0.2967 0.5794 1.4760
CSMLF [27] 0.4441 0.3369 0.2815 0.2408 0.1576 0.4018 0.9378
CSMLF-FT [27] 0.5998 0.4583 0.3869 0.3433 0.2475 0.5018 0.7555
SVM-DBOW [6] 0.7787 0.6835 0.6023 0.5305 0.3797 0.6992 2.2722
SVM-DCONC [6] 0.7547 0.6711 0.5970 0.5308 0.3643 0.6746 2.2222
TrTr-CMR [28] 0.8270 0.6994 0.6002 0.5199 0.3803 0.7220 2.2728
TextGCN [3] 0.7680 0.6892 0.6261 0.5786 0.4009 0.7314 2.8595
FE-LWS 0.8349 0.7589 0.6952 0.6369 0.4554 0.7767 2.9230
TABLE XI: Experimental Results of Baseline Methods and Different Searching Techniques on UCM Dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
R-BOW [22] 0.4107 0.2249 0.1452 0.1095 0.1098 0.3439 0.3071
L-FV [22] 0.5897 0.4668 0.4080 0.3683 0.2698 0.5595 1.8438
CSMLF [27] 0.3874 0.2145 0.1253 0.0915 0.0954 0.3599 0.3703
CSMLF-FT [27] 0.3671 0.1485 0.0763 0.0505 0.0944 0.2986 0.1351
SVM-DBOW [6] 0.7635 0.6664 0.5869 0.5195 0.3654 0.6801 2.7142
SVM-DCONC [6] 0.7653 0.6947 0.6417 0.5942 0.3702 0.6877 2.9228
TrTr-CMR [28] 0.8156 0.7091 0.6220 0.5469 0.3978 0.7442 2.4742
TextGCN [3] 0.8461 0.7844 0.7386 0.6930 0.4868 0.8071 3.4077
FE-LWS 0.8805 0.8305 0.7862 0.7511 0.5014 0.8497 3.6806
TABLE XII: Experimental Results of Baseline Methods and Different Searching Techniques on RSICD Dataset
METHOD BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
R-BOW [22] 0.4401 0.2383 0.1514 0.1041 0.1684 0.3605 0.4667
L-FV [22] 0.4342 0.2453 0.1634 0.1175 0.1712 0.3818 0.6531
CSMLF [27] 0.5759 0.3859 0.2832 0.2217 0.2128 0.4455 0.5297
CSMLF-FT [27] 0.5106 0.2911 0.1903 0.1352 0.1693 0.3789 0.3388
SVM-DBOW [6] 0.6112 0.4277 0.3153 0.2411 0.2303 0.4588 0.6825
SVM-DCONC [6] 0.5999 0.4347 0.3355 0.2689 0.2299 0.4577 0.6854
TrTr-CMR [28] 0.6201 0.3937 0.2671 0.1932 0.2399 0.4895 0.7518
TextGCN [3] 0.6513 0.4819 0.3747 0.3085 0.2752 0.4804 0.8266
FE-LWS 0.6796 0.5124 0.4073 0.3365 0.2944 0.5248 0.9423
Tables XII, XII and XII compares various baseline methods. R-BOW and L-FV [22] use handcrafted sentence representations with RNN and LSTM decoders, respectively. CSMLF and its fine-tuned variant [27] extend this for multi-sentence captioning. SVM-DBOW and SVM-DCONC [6] apply SVMs with bag-of-words and concatenated sentence features. TrTr-CMR [28] combines a Swin Transformer encoder with a cross-modal transformer decoder. TextGCN [3] aligns the embeddings of the decoder with frozen TextGCN output. Our model, FE-LWS, combines dual CNN encoders with stacked GRUs and local weighted aggregation to enhance semantic representation. This method improves performance by effectively integrating the encoding and decoding strategies. The results demonstrate that the proposed FE-LWS model achieves substantially superior performance compared to the baseline methods.
The qualitative evaluation is provided in the supplementary material, where the captions are evaluated by a human annotator [^1] The annotator has several years of experience in RSIC..
Figure 3: Examples of RS Image Captioning by Different Methods
IV-C Visual Interpretation
We present qualitative results from the test sets of three datasets in Figure 3, comparing the proposed model (FE-LWS) with ground-truth captions (Gold) and conventional encoder-decoder models (ConvNext and ResNet). In the first two examples (Figures 3a and 3b), the proposed model correctly classifies the images, while ConvNext and ResNet fail. In Figure 3a, ConvNext and ResNet misclassify the image as industrial area, whereas the proposed model accurately identifies it. In Figure 3b, ConvNext misclassifies the image as resort and ResNet as school, while the proposed model generates a meaningful and accurate caption. In the next two examples (Figures 3c and 3d), the proposed model enhances the quality of the captions by incorporating more detailed information. In Figure 3c, all models detect bridge and river, while the proposed model also captures trees. In Figure 3d, while all models detect park, the proposed model further identifies lake and road.
V Conclusion
Existing RSIC methods focus primarily on decoder enhancements, while giving limited attention to optimizing feature extraction during encoding. Moreover, most approaches rely on a single encoder and decoder, which restricts the model’s ability to handle the complex spatial and semantic characteristics of remote sensing images. To address these limitations, we propose a framework that combines features from two complementary CNN encoders and employs a stacked GRU-based decoder with a locally weighted aggregation strategy. This combination enables a more effective integration of multilevel semantic information, resulting in captions that are contextually rich and semantically accurate. The experimental results across multiple datasets demonstrate the superior performance of the proposed method compared to existing baselines. Fusion-based encoding improves spatial detail extraction, while weighted decoder stacking improves semantic coherence. Although the model shows robust performance, challenges may still arise when dealing with previously unseen or underrepresented image classes. Future research will explore directions to further enhance the generalization and robustness of the model without compromising efficiency. These advances are expected to improve the adaptability and reliability of RSIC in broader applications.
VI Datasets Used
The specifications of the three datasets used in our model are outlined below.
  • SYDNEY: SYDNEY [16] dataset consists of $613$ images, with 497 allocated for training, $58$ for testing, and $58$ for validation. It is derived from the Sydney dataset [29] and has been carefully curated to include seven distinct categories: airportindustrialmeadowoceanresidentialriver, and runway, through thoughtful cropping and selection.
  • UCM: UCM [16] dataset consists of $2,100$ images, with $1,680$ allocated for training, $210$ for testing, and 212121210 for validation. It is derived from the UC Merced Land Use dataset [30], which features $21$ land use image classes, each containing $100$ images. These classes include agricultureairportbaseball diamondbeachbuildingschaparraldenseresidentialforestfreewaygolfcourseharbourintersectionmediumresidentialmobilehomeparkoverpassparkingriverrunwaysparseresidentialstoragetanks and tenniscourt.
  • RSICD: RSICD [22] dataset includes a substantial collection of $10921$ images, with $8034$ designated for training, $1093$ for testing and $1094$ for validation. Sourced from a variety of platforms such as Google Earth [31], Baidu Map, MapABC, and Tianditu, this dataset encompasses $31$ distinct image classes, including airportbarelandbaseballfieldbeachbridgecenterchurchcommercialdenseresidentialdesertfarmlandforestindustrialmeadowmediumresidentialmountainparkparkingplaygroundpondportrailwaystationresortriverschoolsparseresidentialsquarestadiumstoragetanks, and viaduct.
VII Performance Metrices
The evaluation metrics used to assess our model’s performance are described below.
  • BLEU: Bilingual Evaluation Understudy [23] measures n-gram overlap between generated and reference captions. Calculate the geometric mean of the precision of n grams (BLEU-1 to BLEU-4) and apply a brevity penalty to discourage short outputs.
  • METEOR: Metric for Evaluation of Translation with Explicit Ordering [24] incorporates stemming, synonymy, and word order. It balances precision and recall with an emphasis on alignment and content accuracy.
  • ROUGE: Recall-Oriented Understudy for Gisting Evaluation [25] is a recall-based metric that measures n-gram overlap. We use ROUGE-L, which focuses on the longest common subsequence (LCS).
  • CIDEr: Consensus-based Image Description Evaluation [26] evaluates the similarity between generated and human-written captions by measuring consensus across multiple references, highlighting the relevance and agreement of the content.
VIII Subjective Evaluation
TABLE XIII: Subjective Evaluation of Different CNNs on Three Datasets (in %)
Method SYDNEY UCM RSICD
R W U R W U R W U
ConvNext 89.66 1.72 8.62 89.04 4.29 6.67 82.25 5.31 12.44
ResNet 87.93 3.45 8.62 88.09 5.24 6.67 81.43 5.03 13.54
FE-LWS 94.83 3.45 1.72 92.86 5.24 1.90 83.72 6.95 9.33
Unlike classification, multiple captions can describe the same image with similar effectiveness. Hence, relying solely on the test captions (five for all three data sets) to evaluate the predicted captions is not justified [22, 4]. For this reason, we performed a subjective evaluation. In this process, we assigned a human annotator with expertise in this field to assess the quality of the captions generated. This evaluation provides insights into caption quality from the perspective of human psychology. In our work, we used three labels to assess the quality of the generated captions, as discussed below.
  • Related: The generated caption accurately describes the input image.
  • Partially Related: The generated caption describes the input image but contains major errors or flaws.
  • Unrelated: The generated caption is completely incorrect and does not describe the input image in any meaningful way.
Table XIII presents the subjective evaluation results, comparing our model (FE-LWS) with conventional encoder-decoder models, where ConvNext and ResNet are used as encoders, respectively. The results indicate that our model outperforms other models.